# Albanese fibrations of surfaces with low slope

### v2

@@ Abstract.
surface of general type, Albanese fibration
@@ 1. Introduction
We work over the complex number $\mathbb{C}$ throughout this paper. Let $S$ be a minimal irregular surface of general type, and $a:\,S\to\mathrm{Alb}(S)$ its Albanese map. When the image $a(S)$ is a curve, the Albanese map induces a fibration, which we call the Albanese fibration of $S$:
@@ 6. $(vi)$
(1).  Let $f:\,S\to C$ be a hyperelliptic Albanese fibration of genus $g$ with $K_{S}^{2}\leq 4\chi(\mathcal{O}_{S})$. According to the characterization in Theorem 1.2, there should be a better upper bound on $g$. In fact, by [LL24, Theorem 4.1] it is proved that $g\leq\frac{4\chi(\mathcal{O}_{S})+4}{2+\big{(}4\chi(\mathcal{O}_{S})-K_{S}^{2}\big{)}}\leq 2\chi(\mathcal{O}_{S})+2$; see also [Ish05, Theorem 0.1] for the case when $K_{S}^{2}=4\chi(\mathcal{O}_{S})$.
@@
(3). Professor Z. Jiang kindly informed us that in a preprint [JL24] joint with Lin they can also obtain an explicit upper bound on the genus of the Albanese fibration when $p_{g}(S)=q(S)=1$ and $K_{S}^{2}=4$ or $5$.
@@
Can we classify the Albanese fibrations with the equality in (1.3) when $\chi(\mathcal{O}_{S})\leq 4$? We even don’t know whether such Albanese fibrations are bielliptic or not?
@@ 2.1. The surface fibrations
Let $f:S\rightarrow C$ be a fibration of curves of genus $g\geq 2$, i.e. $f$ is a proper morphism from a smooth projective surface $S$ onto a smooth projective curve $C$ with connected fibers over the complex number, and the general fiber is a smooth projective curve of genus $g$. The fibration $f$ is called relatively minimal if there is no $(-1)$-curve (i.e. a smooth rational curve with self-intersection $-1$) contained in the fiber of $f$. It is called hyperelliptic if its general fiber is hyperelliptic, smooth if all its fibers are smooth, isotrivial if all its smooth fibers are isomorphic to each other, and locally trivial if it is both smooth and isotrivial.
Let $\omega_{S}$ (resp. $K_{S}$) be the canonical sheaf (resp. canonical divisor) of $S$, and let $\omega_{f}=\omega_{S/C}=\omega_{S}\otimes f^{*}\omega_{C}^{\vee}$ (resp. $K_{f}=K_{S/C}=K_{S}-f^{*}K_{C}$) the relative canonical sheaf (resp. the relative canonical divisor) of $f$. Put $p_{g}(S):=h^{0}(S,\omega_{S})$, $q(S):=h^{1}(S,\omega_{S})$, $\chi(\mathcal{O}_{S}):=p_{g}(S)-q(S)+1$, and let $\chi_{top}(S)$ be the topological Euler characteristic of $S$. The basic invariants of $f$ are:
@@ 2.2. Bielliptic surfaces
In this subsection, we recall some facts about bielliptic surfaces, and refer to [BHPV04, Bea78, Ser90] for more details. Let $Y$ be a bielliptic surface. The Kodaira dimension of $Y$ is zero, and the invariants of $Y$ is as follows.
@@ 2.3. Xiao’s technique and the second multiplication map
Let now $f:\,S\to B$ be a surface fibration of genus $g\geq 2$. Then $f$ is called a double cover fibration of type $(g,\gamma)$, if there is a rational two-to-one map $\pi:\,S\to Y$, where $h:\,Y\to B$ is a surface fibration of genus $\gamma$ and the following diagram commutes:
@@ 3.1. The bielliptic Albanese fibration
Now return back to the fibration $\tilde{f}:\,\tilde{S}\to\tilde{E}$ obtained by unramified base as above. Note that the general fiber $\widetilde{F}$ of $\widetilde{f}$ are the same as a general fiber of $F$ of $f$. Compare the H-N filtration of $\mathcal{E}=f_{*}\omega_{S/E}$ and ~$\widetilde{\mathcal{E}}:=\widetilde{f}_{*}\omega_{\widetilde{S}/{\widetilde{E}}}$, we see that $\iota_{n-1}:F\rightarrow\Gamma\subset\mathbb{P}^{g-2}$ is the same as $\widetilde{\iota}_{n-1}:\widetilde{F}\rightarrow\widetilde{\Gamma}\subset\mathbb{P}^{g-2}$. In particular, the map $\varphi_{n-1}:S\rightarrow\mathbb{P}_{E}(\mathcal{A})$ defined by the linear system $\Lambda(\mathcal{A})=\Lambda(\mathcal{E}_{n-1})$ factors through $Y$ and we have the following commutative diagram
@@ 3.2.1. The case when $\chi(\mathcal{O}_{S})=4$ and $\omega_{S}^{2}=16$
We will prove Section 3 by contradiction. Suppose (3.2) does not hold, we can assume $14\leq g\leq 15$ by Section 3.2. Note $\omega_{f}^{2}=K_{S}^{2}=16$ since $g(E)=1$. Combining with (2.4), we have
@@
For the subcase $a_{1}=5$ in (case II-1), we have either $a_{2}=3,~{}6\leq a_{3}\leq 7$ or $a_{2}=4,~{}5\leq a_{3}\leq 6$. Note $\frac{g}{2}\geq 7=r_{2}-1\geq r_{1}-1$, by Section 2.3, we have $d_{1}\geq 3(a_{1}-1)$ and $d_{2}\geq 3(a_{1}+a_{2}-1)$. Then by Section 2.3, we have
@@
We will prove Section 3 by contradiction. Suppose (3.2) does not hold, by Section 3.2, we can assume $14\leq a_{1}+a_{2}+a_{3}+a_{4}\leq 15$. Then we get $a_{1}=2,a_{2}=3,a_{3}=4$ and $5\leq a_{4}\leq 6$. By (2.4), we have
@@ 3.2.2. The case when $\chi(\mathcal{O}_{S})=3$ and $\omega_{S}^{2}=12$
We will prove Section 3 by contradiction. Suppose (3.2) does not hold, we can assume $11\leq a_{1}+a_{2}+a_{3}+a_{4}\leq 15$ by Section 3.2. Hence we get $a_{1}\leq 4$. By (2.4), we have
@@ 3.2.3. The case when $\chi(\mathcal{O}_{S})=2$ and $\omega_{f}^{2}=8$
We will prove Section 3 by contradiction. Suppose (3.2) does not hold, we can assume $8\leq a_{1}+a_{2}+a_{3}+a_{4}\leq 15$ by Section 3.2. Hence we get $a_{1}\leq 4$. By (2.4), we have
@@ 3.3. The case when $K_{S}^{2}<4\chi(\mathcal{O}_{S})$
By [LZ18], one has $g\leq 15$. Suppose the conclusion does not hold, then
@@
Moreover, by [BZ00], we know $f_{*}\omega_{S/E}$ is ample. Then we can exclude case (i) and case (ii) using a case-by-case study similarly as in subsubsection 3.2.2 and subsubsection 3.2.1 respectively.
@@
We will prove Section 3 by contradiction. Suppose $\chi(\mathcal{O}_{S})=3,K_{S}^{2}=11,g=11$, Note $\omega_{f}^{2}=K_{S}^{2}=11$ since $g(E)=1$. Combining with (2.4), we have
@@ 4. Examples
In other word, we obtain a sequence of surfaces $S_{n}$ of general type with an Albanese fibration of $g_{n}$ increasing quadratic with $\chi(\mathcal{O}_{S_{n}})$, i.e.,
@@
is base-point-free. According Bertini’s theorem, one can find a smooth divisor
@@
The construction is stimulated by [LL24, Theorem 5.1]. For any integer $m\geq 1$, let $G_{m}=\mathbb{Z}/{m\mathbb{Z}}$ be the cyclic group of order $m$. Suppose that $\sigma\in G_{m}$ acts on $\mathbb{P}^{1}$ by $\sigma(t)=\xi t$ (where $\xi$ is any $m$-th primitive unit root), and $\sigma\in G_{m}$ acts on some elliptic curve $E_{0}$ by translation $\sigma(x)=x+p$, where $p$ is a torsion point of order $m$. Let $P:=(E_{0}\times\mathbb{P}^{1})/{G_{m}}$, where $G_{m}$ acts on $E_{0}\times\mathbb{P}^{1}$ by the diagonal action. Then we have the following commutative diagram
@@ 3. $(3)$
(1) This is in fact contained in the proof of Section 3. Indeed, since $g\geq 5$ and $E$ is an elliptic curve, if follows that the slope of $f$ is $\lambda_{f}=\frac{K_{S}^{2}}{\chi(\mathcal{O}_{S})}=4$. Hence our conclusion follows from [Bar01, Theorem 2.1].
(2) The proof is similar to that of Section 3. Using the same notations as in the proof of Section 3. By (3.5) and (3.6), we have
@@
Thus the upper bound follows according to the description of $\beta$ in each type bielliptic surfaces in Table 2.1 in Section 2.2.

### v3

@@ Abstract.
surface of general type, Albanese fibration, slope
@@
This work is supported by National Natural Science Foundation of China, Shanghai Pilot Program for Basic Research, Fundamental Research Funds for central Universities, Science and Technology Commission of Shanghai Municipality (No. 22DZ2229014) and Natural Science Foundation of Shandong Province (No. ZR2023QA00).
@@ 1. Introduction
We work over the complex number $\mathbb{C}$ throughout this paper. Let $S$ be a minimal irregular surface of general type, and $a:\,S\to\mathrm{Alb}(S)$ be its Albanese map. When the image $a(S)$ is a curve, the Albanese map induces a fibration, which we call the Albanese fibration of $S$:
@@ 6. $(vi)$
(1).  Let $f:\,S\to C$ be a hyperelliptic Albanese fibration of genus $g$ with $K_{S}^{2}\leq 4\chi(\mathcal{O}_{S})$. According to the characterization in Theorem 1.2, there should be a better upper bound on $g$. In fact, in [LL24, Theorem 3.1] it is proved that $g\leq\frac{4\chi(\mathcal{O}_{S})+4}{2+\big{(}4\chi(\mathcal{O}_{S})-K_{S}^{2}\big{)}}\leq 2\chi(\mathcal{O}_{S})+2$; see also [Ish05, Theorem 0.1] for the case when $K_{S}^{2}=4\chi(\mathcal{O}_{S})$.
@@
(3). Professor Z. Jiang kindly informed us that in a preprint [JL24] joint with Lin they can also obtain some explicit upper bounds on the genus of the Albanese fibration when $p_{g}(S)=q(S)=1$ and $K_{S}^{2}=4$ or $5$.
@@
Can we classify the Albanese fibrations with the equality in (1.3) when $\chi(\mathcal{O}_{S})\leq 4$? We even don’t know whether such Albanese fibrations are bielliptic or not.
@@ 2.1. The surface fibrations
Let $f:S\rightarrow C$ be a fibration of curves of genus $g\geq 2$, i.e. $f$ is a proper morphism from a smooth projective surface $S$ onto a smooth projective curve $C$ with connected fibers over the complex number, and the general fiber is a smooth projective curve of genus $g$. The fibration $f$ is called relatively minimal if there is no $(-1)$-curve (i.e. a smooth rational curve with self-intersection $-1$) contained in the fibers of $f$. It is called hyperelliptic if its general fiber is hyperelliptic, smooth if all its fibers are smooth, isotrivial if all its smooth fibers are isomorphic to each other, and locally trivial if it is both smooth and isotrivial.
Let $\omega_{S}$ (resp. $K_{S}$) be the canonical sheaf (resp. canonical divisor) of $S$, and let $\omega_{f}=\omega_{S/C}=\omega_{S}\otimes f^{*}\omega_{C}^{\vee}$ (resp. $K_{f}=K_{S/C}=K_{S}-f^{*}K_{C}$) be the relative canonical sheaf (resp. the relative canonical divisor) of $f$. Put $p_{g}(S):=h^{0}(S,\omega_{S})$, $q(S):=h^{1}(S,\omega_{S})$, $\chi(\mathcal{O}_{S}):=p_{g}(S)-q(S)+1$, and let $\chi_{top}(S)$ be the topological Euler characteristic of $S$. The basic invariants of $f$ are:
@@ 2.2. Bielliptic surfaces
In this subsection, we recall some facts about bielliptic surfaces, and refer to [BHPV04, Bea78, Ser90] for more details. Let $Y$ be a bielliptic surface. The Kodaira dimension of $Y$ is zero, and the invariants of $Y$ are as follows.
@@ 2.3. Xiao’s technique and the second multiplication map
Let now $f:\,S\to B$ be a surface fibration of genus $g\geq 2$. Then $f$ is called a double cover fibration of type $(g,\gamma)$, if there is a rational two-to-one map $\pi:\,S\dashrightarrow Y$, where $h:\,Y\to B$ is a surface fibration of genus $\gamma$ and the following diagram commutes:
@@ 3.1. The bielliptic Albanese fibration
Now return back to the fibration $\tilde{f}:\,\tilde{S}\to\tilde{E}$ obtained by unramified base change as above. Note that the general fiber $\widetilde{F}$ of $\widetilde{f}$ is the same as a general fiber $F$ of $f$. Compare the H-N filtration of $\mathcal{E}=f_{*}\omega_{S/E}$ and ~$\widetilde{\mathcal{E}}:=\widetilde{f}_{*}\omega_{\widetilde{S}/{\widetilde{E}}}$, we see that $\iota_{n-1}:F\rightarrow\Gamma\subset\mathbb{P}^{g-2}$ is the same as $\widetilde{\iota}_{n-1}:\widetilde{F}\rightarrow\widetilde{\Gamma}\subset\mathbb{P}^{g-2}$. In particular, the map $\varphi_{n-1}:S\rightarrow\mathbb{P}_{E}(\mathcal{A})$ defined by the linear system $\Lambda(\mathcal{A})=\Lambda(\mathcal{E}_{n-1})$ factors through $Y$ and we have the following commutative diagram
@@ 3.2.1. The case when $\chi(\mathcal{O}_{S})=4$ and $\omega_{S}^{2}=16$
We will prove Section 3 by contradiction. Suppose (3.2) does not hold. We can assume that $14\leq g\leq 15$ by Section 3.2. Note $\omega_{f}^{2}=K_{S}^{2}=16$ since $g(E)=1$. Combined with (2.4), we get
@@
For the subcase $a_{1}=5$ in (case II-1), we have either $a_{2}=3,~{}6\leq a_{3}\leq 7$ or $a_{2}=4,~{}5\leq a_{3}\leq 6$. Note $\frac{g}{2}\geq 7=r_{2}-1\geq r_{1}-1$. By Section 2.3, we have $d_{1}\geq 3(a_{1}-1)$ and $d_{2}\geq 3(a_{1}+a_{2}-1)$. Then by Section 2.3, we have
@@
We will prove Section 3 by contradiction. Suppose (3.2) does not hold. We can assume that $14\leq a_{1}+a_{2}+a_{3}+a_{4}\leq 15$ by Section 3.2. Then we get $a_{1}=2,a_{2}=3,a_{3}=4$ and $5\leq a_{4}\leq 6$. By (2.4), we have
@@ 3.2.2. The case when $\chi(\mathcal{O}_{S})=3$ and $\omega_{S}^{2}=12$
We will prove Section 3 by contradiction. Suppose (3.2) does not hold. We can assume that $11\leq a_{1}+a_{2}+a_{3}+a_{4}\leq 15$ by Section 3.2. Hence we get $a_{1}\leq 4$. By (2.4), we have
@@ 3.2.3. The case when $\chi(\mathcal{O}_{S})=2$ and $\omega_{f}^{2}=8$
We will prove Section 3 by contradiction. Suppose (3.2) does not hold. We can assume that $8\leq a_{1}+a_{2}+a_{3}+a_{4}\leq 15$ by Section 3.2. Hence we get $a_{1}\leq 4$. By (2.4), we have
@@ 3.3. The case when $K_{S}^{2}<4\chi(\mathcal{O}_{S})$
By [LZ18], one has $g\leq 15$. Suppose the conclusion does not hold. We have
@@
Moreover, by [BZ00], we know that $f_{*}\omega_{S/E}$ is ample. Then we can exclude case (i) and case (ii) using a case-by-case study similarly as in subsubsection 3.2.2 and subsubsection 3.2.1 respectively.
@@
We will prove Section 3 by contradiction. Suppose $\chi(\mathcal{O}_{S})=3,K_{S}^{2}=11,g=11$. Note $\omega_{f}^{2}=K_{S}^{2}=11$ since $g(E)=1$. Combined with (2.4), we get
@@ 4. Examples
In other words, we obtain a sequence of surfaces $S_{n}$ of general type with an Albanese fibration of $g_{n}$ increasing quadratically with $\chi(\mathcal{O}_{S_{n}})$, i.e.,
@@
is base-point-free. According to Bertini’s theorem, one can find a smooth divisor
@@
The construction is stimulated by [LL24, Example 5.1]. For any integer $m\geq 1$, let $G_{m}=\mathbb{Z}/{m\mathbb{Z}}$ be the cyclic group of order $m$. Suppose that $\sigma\in G_{m}$ acts on $\mathbb{P}^{1}$ by $\sigma(t)=\xi t$ (where $\xi$ is any $m$-th primitive unit root), and $\sigma\in G_{m}$ acts on some elliptic curve $E_{0}$ by translation $\sigma(y)=y+p$, where $p$ is a torsion point of order $m$. Let $P:=(E_{0}\times\mathbb{P}^{1})/{G_{m}}$, where $G_{m}$ acts on $E_{0}\times\mathbb{P}^{1}$ by the diagonal action. Then we have the following commutative diagram
@@ 3. $(3)$
(1) This is in fact contained in the proof of Section 3. Indeed, since $g\geq 5$ and $E$ is an elliptic curve, it follows that the slope of $f$ is $\lambda_{f}=\frac{K_{S}^{2}}{\chi(\mathcal{O}_{S})}=4$. Hence our conclusion follows from [Bar01, Theorem 2.1].
(2) The proof is similar to that of Section 3. We use the same notations as in the proof of Section 3. By (3.5) and (3.6), we have
@@
Thus the upper bound follows from the description of $\beta$ in each type bielliptic surfaces in Table 2.1 in Section 2.2.
@@
Data availability  The author declares that the manuscript has no associated data.
Declarations
On behalf of all authors, the corresponding author states that there is no conflict of interest.